Can someone out there help me solve this problem?
At what rate would a 120=pound person burn calories while bicycling at 12 mi/h?
(For a 150-pound person bicycling burns 410 calories)
Thanks,
Sher

Answer by checkley71(8403)

150/120=410/X OR 150X=120*410 OR 150X=49,200 OR X=49,200/150 OR X=328 CALORIES